# SpineBench: A Clinically Salient, Level-Aware Benchmark Powered by the SpineMed-450k Corpus

### v2

@@ Abstract
footnotetext: $\clubsuit$ Equal Contributions. 🖂Corresponding author: suntiansheng-@163.com; cfshan@nju.edu.cn
@@ 1 Introduction
To characterize the state of the field, we evaluate a dozen of contemporary large vision–language models (LVLMs) (OpenAI, 2025a; b; Hurst et al., 2024; Google, 2025a; b; Sellergren et al., 2025; xAI, 2025; Anthropic, 2025; Bai et al., 2025; Hong et al., 2025; Wang et al., 2025a), both general-purpose and medical. Our evaluation reveals significant weaknesses in fine-grained, level-specific diagnosis and open-ended clinical reasoning, particularly in the handling of complex multi-image tasks. Building on these insights, we introduce a fine-tuned spine model SpineGPT trained on SpineMed-450k that delivers consistent improvements on SpineBench as shown in Figure 1. Clinicians assess exemplar outputs for decision relevance, underscoring the practical value of targeted, evidence-linked instruction data. Our contributions are as follows:
@@ 2.1 Data Collection
To build a complete and comprehensive dataset for spinal diagnosis and treatment, we collected data from a variety of sources (Chen et al., 2024a; Wei & Hwei, 2024; Wu et al., 2025; Chen et al., 2024b). Existing general-purpose large vision-language models (Hurst et al., 2024; Google, 2025a; b; Deng et al., 2023; Ullah et al., 2024; AlSaad et al., 2024) and even medical large language models (Li et al., 2023; Wang et al., 2025b; Wu et al., 2024; Lin et al., 2025; Lu et al., 2024; Niu et al., 2025; Nath et al., 2025a; Seyfioglu et al., 2024; Lai et al., 2025; Xu et al., 2025) are trained on generic medical data (Chen et al., 2024a; b; Xie et al., 2024a), which often lacks the high-quality, specialized data needed for orthopedics (Deng et al., 2023; Ullah et al., 2024). To train an effective large model for spinal care, we first compiled a high-quality, general orthopedic dataset covering multiple domains, including Spine Surgery, Foot and Ankle Surgery, Orthopedic Trauma, and Hand and Upper Extremity Surgery.
@@ 2.2 Dataset Curation
To accurately extract comprehensive information from academic sources, we employed PaddleOCR (Du et al., 2020) to parse PDF documents and images from textbooks and literature. The output, containing both recognized text and layout analysis, was exported into Markdown format. This approach effectively preserved the structural integrity of the documents, including tables, figure placements, and overall layout. Furthermore, to ensure a precise mapping between figures, their captions, and corresponding contextual descriptions in the text, we developed a novel algorithm termed Picture Context Matching. The technical details of this algorithm are elaborated in the Appendix F.
@@
This stage focused on processing data sourced from a collection of clinical records in hospitals. We first performed a rigorous de-identification process, removing all sensitive and personally identifiable information (PII), such as patient IDs and physical examination details under HIPAA. We also filtered out irrelevant images, such as post-operative photos and non-diagnostic tables. Subsequently, a Expert LLM model was utilized to conduct a fine-grained classification of the data, ensuring the dataset’s purity by excluding non-orthopedic cases. As shown in Figure 2, the orthopedic domain was categorized into 7 classes, with the spine sub-domain further divided into 14 distinct classes. A detailed statistical overview of the dataset distribution across these categories is presented in Figure 4.
@@
In close collaboration with medical experts, we designed a comprehensive annotation schema to generate high-quality, multi-task training data. The annotation process was tailored to the data source: (1) From External Knowledge Sources (e.g., Textbooks): We generated bilingual (Chinese and English) and multimodal (text and image-based) questions in both multiple-choice and open-ended formats using Expert VLM Model with carefully designed prompts. (2) From Opened-spine Datasets: We processed two open-source spinal datasets, Spark and Verse, to generate multi-turn question-and-answer dialogues that simulate doctor-patient interactions. These datasets consist mainly of unimodal 3D image slices (CT and MRI). To ensure consistency, we standardized the inputs by adaptively sampling 25 slices per case under clinical expert supervision. From this, we created over 300 simulated consultations to train models in their conversational abilities within spinal scenarios. (3) From Real Clinical Records: We created multiple-choice questions, multi-turn conversational datasets for patient interviews, and comprehensive spinal diagnostic reports via Expert VLM Model. For prompt design, please refer to the Appendix H
@@ 3.3 Data Source and Question Type Diversity
As shown in Table 1, our data derives from six major sources: Literature, Textbooks, Case Reports, hospitals, and others. Textbooks, being the primary knowledge source for physicians, constitute the largest proportion with 377k entries, while hospital data, though valuable, is limited in quantity, with 9,668 data points generated from nearly 1,000 real cases. As presented in Table  2 and Figure 4(c), question types are categorized into pure text QA, multimodal QA, medical consultations, and clinical reports, with multiple-choice questions comprising the largest proportion. For evaluation convenience, our test set includes only multiple-choice and clinical report formats.
@@ 4.1 Benchmark Construction
The SpineBench was constructed by sampling from the SpineMed-450k dataset. Following the original distribution of SpineMed-450k, we sampled 500 multiple-choice questions and 100 medical reports. This subset incorporates 14 spinal sub-diseases and data from multiple sources (see Figure 4(b) for details).
@@ 4.2 Evaluation Metrics
5 Experiments
@@ 5.1 Implementation Details
This study employs the Qwen2.5-VL-7B-Instruct model within a curriculum learning framework for subsequent training phases, aimed at enhancing the model’s applicability and proficiency in the field of orthopedic spine care. The training process is divided into three stages, each integrating distinct datasets and training strategies to progressively strengthen the model’s performance in spinal health.
@@
In this initial stage, we utilized several publicly available medical text datasets, including medical-o1-reasoning-SFT (Chen et al., 2024a), Medical-R1-Distill-Data (Chen et al., 2024a), and MedThoughts-8K (hw hwei, 2025). Additionally, we incorporated a diverse set of 150,000 multimodal instruction fine-tuning samples uniformly sampled from PubMedVision (Chen et al., 2024b). The primary objective during this phase is to develop the model’s foundational capabilities in the medical field and to enhance its performance across various contexts. Subsequently, we trained on data from the SpineMed-450k dataset that pertained to non-spinal categories. Our findings indicate that this non-spinal data significantly improved the model’s performance on the SpineBench benchmark, highlighting the importance of broadening the knowledge base to enhance task-specific performance.
@@
Finally, we conducted further training through multi-turn dialogues, report generation, and datasets comprising long-chain reasoning instructions. The goal of this stage is to develop the model’s advanced language comprehension and generation abilities, particularly in the contexts of dialogue interaction and report creation. All training details are provided in the Appendix D
5.2 Results on SpineBench
@@
The evaluation results in Table 4 reveal severe limitations of current vision-language models (OpenAI, 2025a; Hurst et al., 2024; Google, 2025a) in medical domain applications. Large-scale open-source models perform particularly poorly: despite having 72B parameters, Qwen2.5-VL-72B (Bai et al., 2025) achieves only 79.88% average performance and a mere 63.80 cumulative score on medical report generation, far below practical application requirements. Even the best-performing open-source model GLM-4.5V (Hong et al., 2025) (83.26%) exhibits a nearly 6-point gap compared to the leading proprietary model Gemini-2.5-Pro (89.23%). This gap is more pronounced in medical report generation, where proprietary models exceed 85 points while open-source models struggle to reach 80. Additional medical report results are in the Appendix E.
@@ 5.3 Ablations of SpineGPT
As shown in Table 5, models trained exclusively on large-scale general medical data (row 2) exhibit significant performance degradation (74.95 vs. 65.31) on SpineBench compared to the baseline model (row 1). This demonstrates that models trained on such data are insufficient for specialized spine diagnostics. The incorporation of our carefully curated general orthopedic non-spine data (row 3) yields substantial performance improvements (82.14 vs. 74.95), validating the importance of domain-aligned training data. We incorporate spine-specific training data (row 5), which further enhances model performance (87.89 vs. 81.11) compared to using only general medical and orthopedic data (row 4).
5.4 Human-Expert Agreement Analysis
@@ 6 CONCLUSIONS, LIMITATIONS, AND FUTURE WORK
D  Training Strategy ........................................................................................................................................................................D
E  Performance Comparison ........................................................................................................................................................................E
F  Picture Context Matching Algorithm ........................................................................................................................................................................F
@@ A.1 Ethics Statement
This work adheres to the ICLR Code of Ethics. In this study, no human subjects or animal experimentation was involved. All datasets used, including SpineMed-450K, were sourced in compliance with relevant usage guidelines, ensuring no violation of privacy. We have taken care to avoid any biases or discriminatory outcomes in our research process. No personally identifiable information was used, and no experiments were conducted that could raise privacy or security concerns. We are committed to maintaining transparency and integrity throughout the research process.
@@ A.2 Reproducibility Statement
Our work will be fully reproducible: we will open-source SpineBench, all questions, the code for running the API and open-source models, all model outputs, and the code for scoring the models. In other words, every part of the project will be made available.

### v3

@@ 1 Introduction
To characterize the state of the field, we evaluate a dozen of contemporary large vision–language models (LVLMs) (OpenAI, 2025a; b; Hurst et al., 2024; Google, 2025b; a; Sellergren et al., 2025a; xAI, 2025; Anthropic, 2025; Bai et al., 2025; Hong et al., 2025; Wang et al., 2025a), both general-purpose and medical. Our evaluation reveals significant weaknesses in fine-grained, level-specific diagnosis and open-ended clinical reasoning, particularly in the handling of complex multi-image tasks. Building on these insights, we introduce a fine-tuned spine model SpineGPT trained on SpineMed-450k that delivers consistent improvements on SpineBench as shown in Figure 1. Clinicians assess exemplar outputs for decision relevance, underscoring the practical value of targeted, evidence-linked instruction data. Our contributions are as follows:
@@ 2.1 Data Collection
To build a complete and comprehensive dataset for spinal diagnosis and treatment, we collected data from a variety of sources (Chen et al., 2024a; Wei and Hwei, 2024; Wu et al., 2025; Chen et al., 2024b). Existing general-purpose large vision-language models (Hurst et al., 2024; Google, 2025b; a; Deng et al., 2023; Ullah et al., 2024; AlSaad et al., 2024) and even medical large language models (Li et al., 2023; Wang et al., 2025b; Wu et al., 2024; Lin et al., 2025; Lu et al., 2024; Niu et al., 2025; Nath et al., 2025a; Seyfioglu et al., 2024; Lai et al., 2025; Xu et al., 2025) are trained on generic medical data (Chen et al., 2024a; b; Xie et al., 2024a), which often lacks the high-quality, specialized data needed for orthopedics (Deng et al., 2023; Ullah et al., 2024). To train an effective large model for spinal care, we first compiled a high-quality, general orthopedic dataset covering multiple domains, including Spine Surgery, Foot and Ankle Surgery, Orthopedic Trauma, and Hand and Upper Extremity Surgery.
@@ 2.2 Dataset Curation
To accurately extract comprehensive information from academic sources, we employed PaddleOCR (Du et al., 2020) to parse PDF documents and images from textbooks and literature. The output, containing both recognized text and layout analysis, was exported into Markdown format. This approach effectively preserved the structural integrity of the documents, including tables, figure placements, and overall layout. Furthermore, to ensure a precise mapping between figures, their captions, and corresponding contextual descriptions in the text, we developed a novel algorithm termed Picture Context Matching. Subsequently, we employed GPT-5-mini(OpenAI, 2025a) to conduct a semantic consistency check, rigorously filtering out instances where the visual content did not align with the retrieved context. The technical details of this algorithm are elaborated in the Appendix E.
@@
This stage focused on processing data sourced from a collection of clinical records in hospitals. We first performed a rigorous de-identification process, removing all sensitive and personally identifiable information (PII), such as patient IDs and physical examination details under HIPAA. We also filtered out irrelevant images, such as post-operative photos and non-diagnostic tables. Subsequently, GPT-5-mini(OpenAI, 2025a) was utilized to conduct a fine-grained classification of the data, ensuring the dataset’s purity by excluding non-orthopedic cases. As shown in Figure 2, the orthopedic domain was categorized into 7 classes, with the spine sub-domain further divided into 14 distinct classes. A detailed statistical overview of the dataset distribution across these categories is presented in Figure 4.
@@
In close collaboration with medical experts, we designed a comprehensive annotation schema to generate high-quality, multi-task training data. The annotation process was tailored to the data source: (1) From External Knowledge Sources (e.g., Textbooks): We generated bilingual (Chinese and English) and multimodal (text and image-based) questions in both multiple-choice and open-ended formats using Gemini-2.5-pro(Google, 2025b) with carefully designed prompts. (2) From Opened-spine Datasets: We processed two open-source spinal datasets, Spark and Verse, to generate multi-turn question-and-answer dialogues that simulate doctor-patient interactions. These datasets consist mainly of unimodal 3D image slices (CT and MRI). To ensure consistency, we standardized the inputs by adaptively sampling 25 slices per case under clinical expert supervision. From this, we created over 300 simulated consultations to train models in their conversational abilities within spinal scenarios. (3) From Real Clinical Records: We created multiple-choice questions, multi-turn conversational datasets for patient interviews, and comprehensive spinal diagnostic reports via locally deployed GLM-4.5V(Hong et al., 2025) to ensure data security. For prompt design, please refer to the Appendix H
@@
2.3 Comparison with Existing Benchmarks
Table 1: Comparison of SpineMed-450k with existing spine imaging datasets. While prior datasets focus on specific perception tasks, our work introduces the first multimodal instruction-tuning corpus designed for full-spectrum clinical reasoning.
Dataset Modality Scale Core Task Workflow Coverage Output Format
RSNA LumbarDISC (Richards et al., 2025) MRI 2.6k Patients Classification Specific (Stenosis Grading) Class Labels
BUU-LSPINE (Klinwichit et al., 2023) X-Ray 3.6k Patients Detection Specific (Spondylolisthesis) Coordinates/Labels
VerSe 2020 (Liebl et al., 2021) CT 300 Patients Segmentation Specific (Anatomy) Voxel Masks
Lumbar Spine MRI (van der Graaf et al., 2024) MRI 218 Patients Segmentation Specific (Structure) Voxel Masks
Spark (Tianchi) (Tianchi, 2020) CT, MRI 150 Patients Classification Specific (Disease ID) Class Labels
[…]
As illustrated in Table 1, existing datasets such as VerSe(Liebl et al., 2021) and RSNA(Richards et al., 2025) are predominantly unimodal and designed for low-level perception tasks like segmentation, detection, or simple classification. While these datasets serve as effective tools for specific anatomical localization or binary disease identification, they fundamentally fail to capture the holistic context required for complex clinical decision-making, limiting their utility in training models for high-level diagnostic synthesis. In contrast, SpineMed-450k represents a significant paradigm shift from "Tool AI" to "Collaborator AI". Our dataset distinguishes itself through three key dimensions: 1. Multimodal Synthesis, requiring the integration of X-ray, CT, and MRI to mirror real-world cross-modal validation; 2. Cognitive Depth, supporting level-aware reasoning rather than simple label outputs; 3. Workflow Completeness, covering the full patient journey with grounded instructions for Structured Imaging Findings, AI Diagnosis, Treatment Recommendations, and Risk Assessment. This effectively fills the cognitive gap left by perception-focused datasets.
@@ 3.3 Data Source and Question Type Diversity
As shown in Table 2, our data derives from six major sources: Literature, Textbooks, Case Reports, hospitals, and others. Textbooks, being the primary knowledge source for physicians, constitute the largest proportion with 377k entries, while hospital data, though valuable, is limited in quantity, with 9,668 data points generated from nearly 1,000 real cases. As presented in Table  3 and Figure 4(c), question types are categorized into pure text QA, multimodal QA, medical consultations, and clinical reports, with multiple-choice questions comprising the largest proportion. For evaluation convenience, our test set includes only multiple-choice and clinical report formats.
@@ 4.1 Benchmark Construction
The SpineBench was constructed by sampling from the SpineMed-450k dataset. Following the original distribution of SpineMed-450k, we sampled 500 multiple-choice questions and 100 medical reports. This subset incorporates 14 spinal sub-diseases and data from multiple sources (see Appendix F for details).
@@ 4.2 Evaluation Metrics
5 SpineGPT: A Specialized Clinical Collaborator
In this section, we introduce SpineGPT to rigorously validate the efficacy of SpineMed-450k.
@@ 5.1 Implementation Details
Table 5: Training configurations across different stages.
Configuration Stage-1 Stage-2 Stage-3
Datasets
PubMedVision-150k
orthopedics-230k
MedThoughts-8K
Medical-R1-Distill
medical-o1-reasoning
Spine-open
Spine-choice
Spine-chat
(+reasoning)
Spine-report
(+reasoning)
Learning Rate 1e-5 1e-5 1e-6
Max Length 16,384 16,384 49,152
DeepSpeed Zero2 Zero2 Zero3
Epochs 1 1 1
We employed the Qwen2.5-VL-7B-InstructBai et al. (2025) as our foundational architecture. All training phases were executed on a high-performance computational node equipped with 8 NVIDIA A100 GPUs, leveraging the ms-swift framework for efficient fine-tuning. To balance training efficiency with memory constraints across different stages, we dynamically adjusted the DeepSpeed optimization strategies. As detailed in Table 5, for the initial stages involving shorter sequences, we utilized DeepSpeed Zero2; for the final stage requiring extensive context modeling (up to 49k tokens), we transitioned to DeepSpeed Zero3 offloading. The global batch size was optimized per stage to maximize GPU utilization while maintaining training stability.
5.2 curriculum learning
This study employs a curriculum learning framework for subsequent training phases, aimed at enhancing the model’s applicability and proficiency in the field of orthopedic spine care. The training process is divided into three stages, each integrating distinct datasets and training strategies to progressively strengthen the model’s performance in spinal health.
@@
In this initial stage, we utilized several publicly available medical text datasets, including medical-o1-reasoning-SFT (Chen et al., 2024a), Medical-R1-Distill-Data (Chen et al., 2024a), and MedThoughts-8K (hw-hwei, 2025). Additionally, we incorporated a diverse set of 150,000 multimodal instruction fine-tuning samples uniformly sampled from PubMedVision (Chen et al., 2024b). The primary objective during this phase is to develop the model’s foundational capabilities in the medical field and to enhance its performance across various contexts. Subsequently, we trained on data from the SpineMed-450k dataset that pertained to non-spinal categories. Our findings indicate that this non-spinal data significantly improved the model’s performance on the SpineBench benchmark, highlighting the importance of broadening the knowledge base to enhance task-specific performance.
@@
Finally, we conducted further training through multi-turn dialogues, report generation, and datasets comprising long-chain reasoning instructions. The goal of this stage is to develop the model’s advanced language comprehension and generation abilities, particularly in the contexts of dialogue interaction and report creation.
6 Experiments
6.1 Comparison and Analysis on SpineBench
@@
The evaluation results in Table 6 reveal severe limitations of current vision-language models (OpenAI, 2025a; Hurst et al., 2024; Google, 2025b) in medical domain applications. Large-scale open-source models perform particularly poorly: despite having 72B parameters, Qwen2.5-VL-72B (Bai et al., 2025) achieves only 79.88% average performance and a mere 63.80 cumulative score on medical report generation, far below practical application requirements. Crucially, domain-specific pre-training alone proves insufficient: the medical-specialized Medgemma-27B(Sellergren et al., 2025b) achieves an even lower average of 76.66%, trailing our model by over 10 points despite being nearly 4 times larger. Even the best-performing open-source model GLM-4.5V (Hong et al., 2025) (83.26%) exhibits a nearly 6-point gap compared to the leading proprietary model Gemini-2.5-Pro (89.23%). This gap is more pronounced in medical report generation, where proprietary models exceed 85 points while open-source models struggle to reach 80. Additional medical report results are in the Appendix D.
@@
Efficiency and Clinical Utility.
While we acknowledge that SpineGPT (87.44%) slightly trails the absolute SOTA model, Gemini-2.5-Pro (89.23%), this comparison highlights a remarkable efficiency: our 7B-parameter model uses less than 7% of the parameters compared to Gemini-2.5-Pro (which exceeds 100B parameters), yet achieves $\sim$98% of its performance. Notably, SpineGPT also outperforms other top-tier models such as GPT-4o (81.60%). This efficiency translates directly to clinical utility, as our lightweight 7B model is safe and efficient enough for local deployment within hospital firewalls, ensuring data privacy without reliance on external cloud APIs.
6.2 Human-Expert Agreement Analysis
@@ 6.3 Ablations of SpineGPT
As shown in Table 7, models trained exclusively on large-scale general medical data (row 2) exhibit significant performance degradation (74.95 vs. 65.31) on SpineBench compared to the baseline model (row 1). This demonstrates that models trained on such data are insufficient for specialized spine diagnostics. The incorporation of the non-spine orthopedic subset derived from our SpineMed-450k corpus (row 3) yields substantial performance improvements (82.14 vs. 74.95), validating the importance of domain-aligned training data. Notably, training exclusively on the spine subset (row 4) achieves an impressive average score of 87.07, reaching nearly 99% of the full model’s performance. This confirms that our high-density spinal instruction data is the decisive factor. Furthermore, the full multi-stage curriculum (Row 6) incorporates this data to reach the peak performance of 87.89, a substantial enhancement over the configuration relying solely on general medical and orthopedic priors (row 5, 81.11).
@@ 7 CONCLUSIONS, LIMITATIONS, AND FUTURE WORK
ACKNOWLEDGEMENTS
This work was supported by the National Key Research and Development Program of China under Grant 2022YFC2407200, and in part by Sanyou Medical, and in part by cash and in-kind funding from Nanjing Kunpeng&Ascend Center of Cultivation and industry partner(s).
@@ ACKNOWLEDGEMENTS
D  Performance Comparison ........................................................................................................................................................................D
E  Picture Context Matching Algorithm ........................................................................................................................................................................E
F  Detailed Distribution Analysis of SpineBench ........................................................................................................................................................................F
@@ A.1 Ethics Statement
This work strictly adheres to the ICLR Code of Ethics. To address potential concerns regarding data provenance, patient privacy, and the redistribution of public or copyrighted materials, we explicitly outline our comprehensive compliance protocols below.
Our data collection strictly adheres to institutional, legal, and copyright frameworks. Patient privacy is guaranteed through rigorous de-identification procedures. Furthermore, to prevent the unauthorized redistribution of copyrighted texts or publicly available benchmark datasets, we employ a “metadata pointer” mechanism and strictly adhere to the principles of transformative use. Specifically, we do not redistribute any raw public medical images, copyrighted paragraphs, original figures, or sensitive clinical records. A detailed breakdown of our data sources, their respective scales, and the specific compliance mechanisms employed is summarized in Table 8.
Table 8: Overview of Data Sources, Scale, and Compliance Mechanisms
Category Scale Source Compliance & Redistribution Strategy
Clinical Data 9,918 11 authorized tertiary orthopedics centers (retrospective) Zero re-identification risk via 2-stage de-identification (admin info erasure + secure OCR masking). No Raw Redistribution: Only a representative de-identified subset is released under Controlled Access requiring a Data Use Agreement (DUA).
Open Datasets 304 Public spine imaging benchmarks (VerSe, Spark) No Image Redistribution: Strict adherence to original terms. We release only instruction annotations via a Metadata Pointer mechanism, requiring users to map our scripts to original images downloaded directly from official repositories.
Literature & Reports 68,021 Europe PMC Open Access subset API Access: Gathered via official RESTful API. Strict filtering applied for Open Access licenses that explicitly permit academic text and data mining.
Textbooks & Guidelines 377,415 Renowned orthopedics textbooks & top Spine journals Transformative Use: Extracted solely for factual medical concepts, diagnostic logic, and reasoning chains to guide LLM synthesis. Absolutely no copying or redistribution of original copyrighted text or figures.
Question Banks 1,090 Publicly accessible medical exams Transformative Use: Utilized purely as structural templates for MCQ reasoning. The final output consists entirely of LLM-generated synthetic data, constituting academic fair use.
@@ A.2 Reproducibility Statement
To maximize transparency and guarantee reproducibility while strictly adhering to institutional data governance and copyright laws, we adopt a tiered open-source release strategy. We will fully open-source the SpineBench evaluation framework, all codebase implementations (training, inference, and evaluation scoring), and the complete SpineMed-450k instruction annotations (JSON files comprising QA pairs and reasoning chains). To prevent the unauthorized redistribution of copyrighted source images from public benchmarks (e.g., VerSe, Spark), we do not release the raw image files. Instead, we employ a "Metadata-Pointer" mechanism, providing explicit data indices that allow researchers to map our open-source instructions to the original images acquired directly from their official repositories. Regarding the highly sensitive proprietary hospital data, full open-sourcing is legally prohibited. However, to ensure the verifiability of our claims, we will release a representative, fully de-identified clinical subset under a Controlled Access model governed by a Data Use Agreement (DUA). This released subset, in conjunction with the public data pointers, provides the exact necessary resources to fully reproduce all SpineBench evaluation metrics reported in this paper.